Puget Sound Divers is a company that provides diving services such as underwater ship repairs to clients in the Puget Sound area. The company’s planning budget for May appears below:

  

Puget Sound Divers
Planning Budget
For the Month Ended May 31
  Budgeted diving-hours (q) 250
   Revenue ($440.00q) $ 110,000
  Expenses:
       Wages and salaries ($11,800 + $128.00q) 43,800
       Supplies ($3.00q) 750
       Equipment rental ($2,100 + $22.00q) 7,600
        Insurance ($4,000) 4,000
       Miscellaneous ($530 + $1.48q) 900
  Total expense 57,050
  Net operating income $ 52,950

  

Required:

During May, the company’s activity was actually 240 diving-hours. Complete the following flexible budget for that level of activity. (Round your answers to nearest whole dollar.)

Puget Sound Divers is a company that provides divi
Answer #1
Concepts and reason

Flexible budget: A flexible budget is a budget that changes with change in volume of production. It flexes the amounts specified in the budget with the change in level of activity. It is an alternative to the static budget that remains constant for any change in level of activity.

Fundamentals

Revenue: Revenue is the amount earned or to be earned by a organization by sale of goods or by providing sale of services. It will have a credit balance and it is part of the income statement.

Expenses: Expense is the amount incurred or to be incurred by a organization for earning revenue. It will have a debit balance and it is part of the income statement.

Net income or loss: Net income or loss is the difference of the revenue earned or to be earned and expenses incurred to be incurred. Net income is the positive difference between the revenue and expenses. Net loss is the negative difference between the revenue and expenses.

Variable cost: The variable costs are those costs which are directly related to the manufacturing process. The variable cost is constant on per unit basis. They will vary with to total variable cost. These are generally considered as relevant costs for decision making processes as they are direct costs. Direct material per unit, labor cost per unit and freight are some of the examples of variable cost.

Fixed cost: A fixed cost is a one-time expense to a company. A fixed cost always remains unchanged without any change though the volume of production changes. Generally fixed costs are in huge amount. Fixed cost per unit changes with the increase or decrease in production. The total fixed costs remain unchanged. Examples of fixed costs are depreciation, advertisement, insurance, rent, etc.

 

Calculate the items of flexible budget as shown below:

Revenue=Revenueperdivinghour×Numberofdivinghours=$440×240=$105,600\begin{array}{c}\\{\rm{Revenue = Revenue per diving hour}} \times {\rm{Number of diving hours }}\\\\{\rm{ = \$ 440}} \times 240\\\\{\rm{ = \$ 105,600}}\\\end{array}

Wagesandsalaries=$11,800+($128×240divinghours)=$11,800+$30,720=$42,520\begin{array}{c}\\{\rm{Wages and salaries = \$ 11,800 + (\$ 128}} \times {\rm{240 diving hours)}}\\\\{\rm{ = \$ 11,800 + \$ 30,720}}\\\\{\rm{ = \$ 42,520}}\\\end{array}

Supplies=$3×240=$720\begin{array}{c}\\{\rm{Supplies = \$ 3}} \times 240\\\\{\rm{ = \$ 720}}\\\end{array}

Equipmentrental=$2,100+($22×240)=$2,100+$5,280=$7,380\begin{array}{c}\\{\rm{Equipment rental = \$ 2,100 + }}(\$ 22 \times 240)\\\\{\rm{ = \$ 2,100 + \$ 5,280}}\\\\{\rm{ = \$ 7,380}}\\\end{array}

Miscelleneous=$530+($1.48×240)=$530+355=$885\begin{array}{c}\\{\rm{Miscelleneous = \$ 530 + (\$ 1}}{\rm{.48}} \times 240)\\\\{\rm{ = \$ 530 + 355}}\\\\{\rm{ = \$ 885}}\\\end{array}

Prepare the flexible budget as shown below:

Therefore, the net operating income under the flexible budget is $50,095

Ans:

The net operating income under the flexible budget is $50,095.